3.What if the desired final mixture was changed to yield a 35% mixture containing pure lemon juice?
a.How many gallons of a 20% solution should be mixed with 15 gallons of a 70% solution to yield a final mixture of lemonade containing 35% of pure lemon juice? Write your answer in complete sentences. Show all work.
b.How many gallons of lemonade would this combination yield? Write your answer in complete sentences. Show all work

Answer by stanbon(75887)

a.How many gallons of a 20% solution should be mixed with 15 gallons of a 70% solution to yield a final mixture of lemonade containing 35% of pure lemon juice? Write your answer in complete sentences. Show all work.
------
0.20x + 0.70*15 = 0.35*(x+15)
-----
20x + 70*15 = 35x + 35*15
-----
15x = 35*15
-----
x = 35 gallons (amt of 20% solution needed)
------------
Cheers,
Stan H.
---------------
b.How many gallons of lemonade would this combination yield? Write your answer in complete sentences. Show all work
Ans: x+15 = 35+15 = 50 gallons
-----------
Cheers,
Stan H.
-----------